# Quantum backflow in biased tight-binding systems

### v1

@@ V.2 Infinite chain
We now generalize the procedure from the previous section to a superposition that includes all eigenstates of the infinite chain with positive momentum, with the aim of determining the lower and upper bounds that constrain the probability density flux.
For this system, according to eq. (14), the most general superposition of states with positive momenta is
@@
The flux associated to the wave function $\psi(j,t)$ in eq. (35) can be computed from eq.(17) as a functional of $\phi(k)$:
@@
$\begin{split}\frac{\tau\sqrt{1+\epsilon^{2}}}{\pi\hbar}\int_{-\xi}^{\pi-\xi}\phi(k)\,e^{i\left((j^{\prime}-\frac{1}{2})(k-k^{\prime})+\frac{(E_{k^{\prime}}-E_{k})t^{\prime}}{\hbar}\right)}\sin\left(\frac{k+k^{\prime}}{2}+\xi\right)\,dk^{\prime}=\lambda\phi(k^{\prime}),\end{split}$ (39)
@@
The last equation is a homogeneous Fredholm integral equation that can be solved by direct computation wazwaz2011linear. First we set
@@
In fig. 2a, we illustrate the probability density flux given by Eq. (49) as a function of time for position $j=j^{\prime}=3$ compared to adjacent sites j1,j+1j-1,j+1, taking $t^{\prime}=3$, so the extreme values of the flux occur at site $j=3$ at time $t=3$ as observed. We also show the bounds associated with the corresponding eigenvalues $\lambda_{\pm}$. The insets show that the largest (smallest) flux values indeed coincide with the upper (lower) bounds at the correct positions and times, as expected.
@@ V.3 Periodic chain
We address now the case when the system has periodic boundary conditions. A general state with nonnegative momentum can be written as the linear combination
@@
remembering that $\eta_{2}=\lfloor\frac{\pi-\xi}{2\pi}N\rfloor$ and $\eta_{1}=\lceil-\frac{\xi}{2\pi}N\rceil$. With these sumation limits, we are taking under consideration all the nonnegative possible flux states, as we discussed in the previous section. The coefficients $c_{m}$ satisfy the normalization condition
@@
The derivation of the bounds is analogous to the infinite-chain case; however, instead of finding a function, we seek to determine the sequence $c_{m}$ that extremizes eq. (52) (see the Appendix for details). Here again we have $\max(J(j,t))=\lambda_{+}$ and $\min(J(j,t))=\lambda_{-}$ where
@@
In fig. 2b, we show the flux computed through Eq. (52) for the sequence defined in Eq. (54), together with their respective bounds $\lambda_{\pm}$, using the same values of $j^{\prime}$ and $t^{\prime}$ as in the infinite chain. The main difference with respect to that case lies in the fact that $J(j;j^{\prime},t;t^{\prime})$ does not decay to zero as $t$ approaches infinity; instead, it oscillates with smaller amplitudes, exhibiting intervals of attenuation and amplification, although it does not appear to reach the same magnitude as at $j^{\prime}$ and $t^{\prime}$.
@@
The choice of $\eta_{2}$ for odd $N$ is made in order to avoid the zero-moment state being considered twice in the state of eq. (35). Thus the bounds of the flux given in eq.(53) become
@@
Thus we have that eq. (54) reduces to
@@
Figure 2: The extreme values of the probability density flux as a at $j^{\prime}=3$, and the nearest sites are shown as a function of $t$ taking $t^{\prime}=3$ (dashed red line), for the infinite chain (a) and the periodic chain (b) with $\epsilon=1$ and $N=9$. We compute the flux using the function $\phi_{\pm}(k)$ (sequence $c_{m}^{\pm}$) and the bound associated : $\lambda_{+}$ (blue dashed line) and $\lambda_{-}$ (magenta dashed line).
The effect of the parameter $\epsilon$ on the probability density flux is shown in fig. (3). We observe that the magnitude of the bounds on the flux increases with $\epsilon$. This increase is more pronounced with $\lambda_{+}$, while it is comparatively milder with $\lambda_{-}$. An additional effect of the bias can be observed for the periodic chain, which exhibits larger oscillations of the flux far from the vicinity of $t=t^{\prime}$, and these oscilations also increase with $\epsilon$.
Figure 3: The extreme values of the probability density flux for different $\epsilon$ values at $j^{\prime}=0$, and $t^{\prime}=3$ (dashed red line) as a time function, for the infinite chain (a) and the periodic chain with $N=9$ (b). The dashed lines are the respective bounds, $\lambda_{+}$ and $\lambda_{-}$ for the corresponding values of $\epsilon$.
@@ VI The Bracken-Melloy approach to backflow
In this section we study backflow using the approach proposed by Bracken and Melloy. The original idea behind this procedure was formulated for a wave function composed by the superposition of states with positive momenta, as above, on a continuous, infinite one dimensional system. However, instead of calculating the extreme values of the flux, it aims to estimate the maximum amount of probability transferred from right to left of the origin, in a time interval $T$. This quantity can be defined as: $c_{\text{BM}}=\max\big(P(T)-P(0)\big)$, where $P(t)$ denotes the probability of finding the particle in the region, say, $x<0$, after a time $t$ Bracken_1994. This is equivalent to finding the upper bound of the integral of the flux over a time window [T/2,T/2][-T/2,T/2], at a given point (normally the origin), i.e. $c_{\text{BM}}=\max\big(-\int_{-T/2}^{T/2}J(0,t)\,dt\big)$. It is important to point out that for systems with periodic boundary conditions only the second definition of $c_{\text{BM}}$ makes sense.
It has been shown that for continuous periodic systems, the total probability that flows to the left at the origin is bounded by $c^{\text{cont}}_{\text{ring}}\approx 3.0380c_{\text{BM}}$ ring1.
@@ VI.1 Infinite chain
Here we proceed to derive the eigenvalue problem to maximize the integral $\int_{-T/2}^{T/2}J(0,t)\,dt$ for the infinite chain. First, we write the flux in eq. (37) at the origin as
@@
where we introduced the adimensional parameter $\nu=\frac{\tau T}{\hbar}$. In order to maximize this quantity subject to the constraint given by eq. (36), we consider the functional:
@@
where $\tilde{\Phi}(p)=\lim_{(\Delta x)\rightarrow 0}\Phi((\Delta x)p)$ and $\zeta$ is defined in eq. (20). This is the "biased" version of the eigenvalue problem obtained by Bracken and Melloy Bracken_1994.
Note that the kernel is the same as the one obtained for a backflow problem under a constant force Melloy1; Goussev1 with the difference that the lower integration limit in eq. (62) is shifted due to the effect of the drift.
As occurs with its continuous counterpart, the eigenvalue problem of eq. (60) cannot be computed analytically, so we approach the problem numerically. We base the calculation on the fact that the integration can be approximated as a symmetric finite square matrix acting on a vector. The largest eigenvalue of the matrix is taken as an approximation to $\lambda_{p}$ Bracken_1994; eveson; Penz_2006. Fig. (4) shows the dependence of the largest eigenvalue on the parameter $\nu$ for several values of $\epsilon$. The value of $c_{\text{BM}}$, which bounds the continuous case, is indicated in the figure. We observe that as $\nu$ increases, $\lambda_{p}$ appears to become bounded from below by $c_{\text{BM}}$. Here, $\epsilon$ appears to act only as some kind of scaling factor for each curve, narrowing it towards the $y$-axis without actually affecting qualitatively the behavior of the curves.
Figure 4: Supremum of the minus time-integrated probability flux, $\min(\lambda_{p})$, as a function of $\nu$ for different values of $\epsilon$. It can be observed that $\lambda_{p}$ appears to become bounded below by $c_{\text{BM}}$ as $\nu\rightarrow\infty$. The maximum of $\lambda_{p}$ tends to be larger for values of $\nu$ close to zero, and the effect of $\epsilon$ does not significantly change the qualitative behavior of the curves. We found that the three curves reach a maximum value of $0.0764734$.
@@ VI.2 Periodic chain
In this section we consider the case of a periodic chain. First, as above, we compute the integral of eq. (51) at $j=0$, over the temporal window [T/2,T/2][-T/2,T/2] :
@@
where the indexes $m,n$ run from $\eta_{1}$ to $\eta_{2}$. This eigenvalue equation (65) can be connected to its continuous counterpart ring1 using
@@
We solve numerically the eigenvalue problem of eq. (65) in the same way as we did for the infinite chain case. Our results are shown in figure (5). There we display the maximum eigenvalue $\lambda_{p}$ for several values of $\epsilon$, highlighting $c_{\text{BM}}$ and $c^{\text{cont}}_{\text{ring}}$, increasing the size of the lattice by an order of magnitude in each figure. We note that as the number of sites increases, the figures resembles to some extent those reported by Goussev ring1. The curve for $\epsilon=0$ exhibits the largest value of $\max(\lambda_{p})$, and this value drops as the lattice grows.
@@
Finally, for each given $N$, we denote by $c_{\text{ring}}^{\text{tb}}$ the maximum of the curve $\max(\lambda_{p})$ as a function of $\nu$ when $\epsilon=0$. In fig. (6) we show how $c_{\text{ring}}^{\text{tb}}$ and its corresponding value of $\nu_{c_{\text{ring}}^{\text{tb}}}$ depend on $N$. To establish a clear relationship, we subtract from $c_{\text{ring}}^{\text{tb}}$ the value of $c^{\text{cont}}_{\text{ring}}=0.11681564947322964$ ring1 to which it tends asymptotically as $N^{-1.7}$ approximately. On the other hand, the right panel shows that to a good approximation, $\nu_{c_{\text{ring}}^{\text{tb}}}$ has a dependence of the form $\sim N^{2}$ for large $N$. However, the overall maximum value of $c_{\text{ring}}^{\text{tb}}$ for this system is reached when $N=5$, reaching $\max(c_{\text{ring}}^{\text{tb}})\approx 0.131349787116051$, which is 12% larger than its continuous counterpart $c^{\text{cont}}_{\text{ring}}$.
@@
Figure 5: Maximum of the time negative integrated probability density flux as a function of $\nu$, for $N=10,10^{2}$ and $10^{3}$ sites. The solid gray line corresponds to the result for the continuous version of the system, while the magenta dotted line denotes $c_{\text{BM}}$. We note that as the number of sites increases, the peak value of the curves converges toward $c^{\text{cont}}_{\text{ring}}$ from above, which is in accordance with the scaling law of eq. (67).
@@
Figure 6: In the left panel, we show $c_{\text{ring}}^{\text{tb}}-c_{\text{ring}}^{\text{cont}}$, and in the right panel the value of $\nu_{c_{\text{ring}}^{\text{tb}}}$, both as functions of $N$. Additionally, we included the fits for the corresponding curves.
We also compute the probability flux at $j=0$ as a function of time for the states that maximize the total backflowing probability, i.e. the solutions to the eigenvalue eqns. (60) and (65). We do this by numerically evaluating the expressions of eqns. (37) and (52). In the left panel we took a $\nu=50$ for the infinite chain, because in fig.( 4), that value corresponds to the point where the oscillations of $\max(\lambda_{p})$ come closest to $c_{\text{BM}}$. In the right panel we took $\nu=26.4$ and $N=20$ for the periodic chain, this corresponds to the value of $c_{\text{ring}}^{\text{tb}}$ for this number of sites.
@@
We stress the differences between the backflow discussed in this section and that analyzed in section V. In section V we determine we determine the largest negative flux that can appear in these systems, whereas in the Bracken and Melloy approach, we focus on the maximum probability that flows to the left of the origin in a time interval $T$. This results from a sustained backflowing flux in a time window [T/2,T/2][-T/2,T/2] with a much more modest amplitude.
@@
Figure 7: Probability density flux at $j=0$ as function of $t$, for the infinite chain (a), and the periodic chain with $N=9$ (b), with several values $\epsilon$ using the eigenfunctions appearing in eqns. (60) and (65) respectively. The black dashed lines indicate the limits of the integration time window.
@@ VII Conclusions and outlook
In this work we have described some of the main characteristics of the quantum backflow phenomenon in tight-binding model with complex couplings, both for infinite chains and for periodic chains. These systems display interesting features not only because they broadens the possible experimental approaches in which this effect could be detected and measured, but also from the transport point of view, since it is analogous to classical biased diffusion systems Zimbor; Lu; Liu or to systems subject to a magnetic field GVugalter_2004. Actually, the detection of backflow is still an experimental challenge, since it has not been observed in the laboratory, although some experimental proposals have been formulated Muga; Yearsley1; Barbier2021experimentfriendly. This has motivated research in quantum backflow to focus on finding configurations that facilitate two aspects: the detection of the effect and the measurement of the quantity $\int_{T/2}^{T/2}J(0,t)dt$. Our present work also aims in this direction: we focused on identifying scenarios in which the negative current is most pronounced and therefore, presumably easier to detect. In this sense, it is similar to the approach used by Palmero et al. Muga in their proposal of an experimental realization based on Bose–Einstein condensates. From this perspective, we found that though it breaks the left-right symmetry in the system, the effect of the bias is essentially to enhance the maximum flux amplitude by factor of $\sqrt{1+\epsilon^{2}}$ with respect to the unbiased case. For completeness, we also calculated the upper bounds of the quantity $\int_{-T/2}^{T/2}J(0,t),dt$, and found that they are larger than their continuous counterparts, e.g., $c_{\text{BM}}=0.0384517$ for a particle in an infinite line Bracken_1994 vs $0.0764734$ for the infinite chain, whereas for the periodic chains, this quantity depends on the number of sites, being larger for small lattice sizes and $\epsilon=0$. Therefore, tight-binding systems may prove to be suitable systems to observe backflow experimentally and open possibilities for future applications based on it.

### v2

@@ V.2 Infinite chain
We now address the general case, starting with the infinite chain. In view of eq. (26), the most general wave function composed of states with positive-momentum is given by
@@
The flux associated with the wave function $\psi(j,t)$ in eq. (35) can be computed from eq.(7) as a functional of $\phi(k)$:
@@
$\begin{split}\frac{\tau\sqrt{1+\epsilon^{2}}}{\pi\hbar}\int_{-\xi}^{\pi-\xi}\phi(k)\,e^{i\left((j^{\prime}-\frac{1}{2})(k-k^{\prime})+\frac{(E_{k^{\prime}}-E_{k})t^{\prime}}{\hbar}\right)}\sin\left(\frac{k+k^{\prime}}{2}+\xi\right)\,dk=\lambda\phi(k^{\prime}),\end{split}$ (39)
@@
The last equation is a homogeneous Fredholm integral equation that can be solved by direct computation Wazwaz (2011). First we set
@@
In fig. 2a, we illustrate the probability density flux given by eq. (49) as a function of time for position $j=j^{\prime}=3$, and its vicinity. We take take $t^{\prime}=3$, so the extreme values of the flux occur at site $j=3$ at time $t=3$ as observed. We also show the bounds associated with the corresponding eigenvalues $\lambda_{\pm}$. The insets show that the largest (smallest) flux values indeed coincide with the upper (lower) bounds at the correct positions and times, as expected.
@@ V.3 Periodic chain
The procedure developed for the infinite chain can now be adapted and applied to the periodic-chain case, taking into account, of course, the fact that the spectrum is discrete. We begin by noting that, according to eq. (27), the most general superposition of eigenstates with non-negative momentum can be written as
@@
recalling that $\eta_{2}=\lfloor\frac{\pi-\xi}{2\pi}N\rfloor$ and $\eta_{1}=\lceil-\frac{\xi}{2\pi}N\rceil$. The coefficients $c_{m}$ satisfy the normalization condition
@@
The derivation of the bounds is analogous to the infinite-chain case; however, instead of finding a function, we seek to determine the sequence $c_{m}$ that extremizes eq. (52) (see the Appendix for details). Here again, we have $\max(J(j,t))=\lambda_{+}$ and $\min(J(j,t))=\lambda_{-}$ where
@@
In fig. 2b, we show the fluxes computed through eq. (52) for the sequence defined in eq. (54), together with their respective bounds $\lambda_{\pm}$, using the same values of $j^{\prime}$ and $t^{\prime}$ as in the infinite chain. The main difference with respect to that case lies in the fact that $J(j;j^{\prime},t;t^{\prime})$ does not rapidly decay to zero as $t$ increases; instead, it oscillates with smaller amplitudes, exhibiting intervals of attenuation and amplification, although it does not appear to reach the same magnitude as at $j^{\prime}$ and $t^{\prime}$.
@@
The choice of $\eta_{2}$ for odd $N$ is made in order to avoid the zero-momentum state being considered twice in the wave function defined in eq.(50). Therefore, the bounds of the flux given in eq.(53) become
@@
Thus, we have that eq. (54) reduces to
@@
Figure 2: The extreme values of the probability density flux at $j^{\prime}=3$, and the nearest sites are shown as a function of $t$ taking $t^{\prime}=3$ (dashed red line), for the infinite chain (a) and the periodic chain (b) with $\epsilon=1$ and $N=9$. We compute the flux using the function $\phi_{\pm}(k)$ (sequence $c_{m}^{\pm}$) and the bound associated : $\lambda_{+}$ (blue dashed line) and $\lambda_{-}$ (blue dashed line).
The effect of the parameter $\epsilon$ on the probability density flux is shown in fig. (3). We observe that the magnitude of the bounds on the flux increases with $\epsilon$. This increase is more pronounced with $\lambda_{+}$, while it is comparatively milder with $\lambda_{-}$. An additional effect of the bias can be observed for the periodic chain, which exhibits larger oscillations of the flux far from the vicinity of $t=t^{\prime}$, and these oscillations also increase with $\epsilon$.
Figure 3: The extreme values of the probability density flux for different $\epsilon$ values at $j^{\prime}=0$, and $t^{\prime}=3$ (dashed red line) as a functions of time, for the infinite chain (a) and the periodic chain with $N=9$ (b). The dashed lines are the respective bounds, $\lambda_{+}$ and $\lambda_{-}$ for the corresponding values of $\epsilon$.
@@ VI The Bracken-Melloy approach to backflow
Following the work of Bracken and Melloy Bracken and Melloy (1994), we now focus on the amount of probability flowing backwards during a time interval in our discrete tight-binding system. This description complements the analysis of the previous section. The original formulation of the problem aimed at estimating the maximum probability for a particle to be transferred from right to left across a given point (typically the origin) within a time interval $T$, for a wave function composed of a superposition of positive-momentum states in an infinite one-dimensional continuous system. This quantity is known as the Bracken–Melloy constant and is defined as
$c_{\text{BM}}=\max\big(P(T)-P(0)\big),$ (57)
where $P(T)$ denotes the probability of finding the particle in the region $x<0$, after a time $T$ Bracken and Melloy (1994). This is equivalent to determining the upper bound of the time integral of the probability current over a time window [T/2,T/2][-T/2,\,T/2] at the origin, namely
$c_{\text{BM}}=\max\!\left(-\int_{-T/2}^{T/2}J(0,t)\,dt\right).$ (58)
It is important to point out that, for systems with periodic boundary conditions, only the second definition of $c_{\text{BM}}$ makes sense.
As an extension of Bracken and Melloy’s work, it has been shown that for continuous periodic systems, the total probability that flows to the left at the origin is bounded by $c^{\text{cont}}_{\text{ring}}\approx 3.0380c_{\text{BM}}$ Goussev (2021).
@@ VI.1 Infinite chain
In this section we compute the extrema of the integrated flux for the case of our discrete infinite chain. First we write the flux in eq. (37) at the origin as
@@
where we introduced the dimensionless parameter $\nu=\frac{\tau T}{\hbar}$. In order to maximize this quantity subject to the constraint given by eq. (36), we consider the functional:
@@
where $\tilde{\Phi}(p)=\lim_{(\Delta x)\rightarrow 0}\Phi((\Delta x)p)$ and $\zeta$ is defined in eq. (10). This is the "biased" version of the eigenvalue problem obtained by Bracken and Melloy Bracken and Melloy (1994).
Note that the kernel is the same as the one obtained for a backflow problem under a constant force Melloy and Bracken (1998); Goussev (2019) with the difference that the lower integration limit in eq. (64) is shifted due to the effect of the drift.
As occurs with its continuous counterpart, the eigenvalue problem of eq. (62) cannot be computed analytically, so we approach the problem numerically. We base the calculation on the fact that the integration can be approximated as a symmetric finite square matrix acting on a vector. The largest eigenvalue of the matrix is taken as an approximation to $\lambda_{p}$ Bracken and Melloy (1994); Eveson et al. (2005); Penz et al. (2005). Fig. (4) shows the dependence of the largest eigenvalue on the parameter $\nu$ for several values of $\epsilon$. The value of $c_{\text{BM}}$, which bounds the continuous case, is indicated in the figure. We observe that as $\nu$ increases, $\lambda_{p}$ appears to become bounded from below by $c_{\text{BM}}$. Here, $\epsilon$ appears to act only as some kind of scaling factor for each curve, narrowing it towards the $y$-axis without actually affecting qualitatively the behavior of the curves.
Figure 4: Supremum of the minus time-integrated probability flux, $\min(\lambda_{p})$, as a function of $\nu$ for different values of $\epsilon$. It can be observed that $\lambda_{p}$ appears to become bounded below by $c_{\text{BM}}$ as $\nu\rightarrow\infty$. The maximum of $\lambda_{p}$ tends to be larger for values of $\nu$ close to zero, and the effect of $\epsilon$ does not significantly change the qualitative behavior of the curves. We find that the three curves reach a maximum value of $0.0764734$.
@@ VI.2 Periodic chain
We now turn our attention to the periodic chain. Using the expressions derived in Section V, the integral of eq. (52) at $j=0$ over the temporal window [T/2,T/2][-T/2,T/2] is:
@@
where the indices $m,n$ run from $\eta_{1}$ to $\eta_{2}$. This eigenvalue equation (67) can be connected to its continuous counterpart Goussev (2021) using
@@
We solve numerically the eigenvalue problem of eq. (67) in the same way as we did for the infinite chain case. Our results are shown in figure (5). There we display the maximum eigenvalue $\lambda_{p}$ for several values of $\epsilon$, highlighting $c_{\text{BM}}$ and $c^{\text{cont}}_{\text{ring}}$, increasing the size of the lattice by an order of magnitude in each figure. We note that as the number of sites increases, the figures increasingly resemble those reported for the continuous ring by Goussev Goussev (2021). Specifically, for $\epsilon=0$, the curves display a shape to those obtained for the continuous ring, they have similar features and their maxima tend to $c^{\text{cont}}_{\text{ring}}$, corresponding to the gray line, as $N$ increases.
@@
Finally, for each given $N$, we denote by $c_{\text{ring}}^{\text{tb}}$ the maximum of the curve $\max(\lambda_{p})$ as a function of $\nu$ when $\epsilon=0$. In fig. (6) we show how $c_{\text{ring}}^{\text{tb}}$ and its corresponding value of $\nu_{c_{\text{ring}}^{\text{tb}}}$ depend on $N$. To establish a clear relationship, we subtract from $c_{\text{ring}}^{\text{tb}}$ the value of $c^{\text{cont}}_{\text{ring}}=0.11681564947322964$ Goussev (2021) to which it tends asymptotically as $N^{-1.7}$ approximately. On the other hand, the right panel shows that to a good approximation, $\nu_{c_{\text{ring}}^{\text{tb}}}$ has a dependence of the form $\sim N^{2}$ for large $N$. However, the overall maximum value of $c_{\text{ring}}^{\text{tb}}$ for this system is reached when $N=5$, reaching $\max(c_{\text{ring}}^{\text{tb}})\approx 0.131349787116051$, which is 12% larger than its continuous counterpart $c^{\text{cont}}_{\text{ring}}$.
@@
Figure 5: Maximum of the time negative integrated probability density flux as a function of $\nu$, for $N=10,10^{2}$ and $10^{3}$ sites. The solid gray line corresponds to the bound for the continuous version of the system, while the blue dotted line denotes $c_{\text{BM}}$. We note that as the number of sites increases, the peak value of the curves converges toward $c^{\text{cont}}_{\text{ring}}$.
@@
Figure 6: In the left panel, we show $c_{\text{ring}}^{\text{tb}}-c_{\text{ring}}^{\text{cont}}$, and in the right panel the value of $\nu_{c_{\text{ring}}^{\text{tb}}}$, both as a functions of $N$. Both curves appear to be well described by a power law. We include the best power law fits for each curve.
In fig. (7), we show the probability density flux at the site $j=0$ as a function of time for the states that maximize the integrated (negative) current. This was obtained by numerically evaluating the expressions in eqs. (37) and (52), using the solutions of the eigenvalue equations (62) and (67), respectively, for two values of $\epsilon$. In the left hand of fig. (7), we present the infinite-chain case and set $\nu=50$, which corresponds to the maximum value of $\nu$ used in Fig. (4). In the right hand of fig. (7), we display the instantaneous flux for the periodic chain with $N=20$ sites. In this case, we choose $\nu=26.4$, which corresponds to the value at which the integrated backflow reaches $c_{\text{ring}}^{\text{tb}}$ for this lattice size.
@@
We stress the differences between the backflow discussed in this section and that analyzed in Section V. In Section V we determine the largest negative instantaneous flux that can appear in these systems, whereas in the Bracken and Melloy approach, we focus on the maximum probability that flows to the left of the origin in a time interval $T$. This results from a flux oscillating with a modest amplitude around a negative value during the time window [T/2,T/2][-T/2,T/2] .
@@
Figure 7: Probability density flux at $j=0$ as a function of $t$, for the infinite chain (a), and the periodic chain with $N=20$ (b), with several values of $\epsilon$ using the eigenfunctions appearing in eqns. (62) and (67) respectively. The black dashed lines indicate the limits of the integration time window.
@@ VII Conclusions and outlook
In this work, we describe the key aspects of quantum backflow in biased tight-binding systems, for both infinite and periodic chains. We first focus on the transport properties of the system, in particular, on the significance and physical implications of the complex hopping parameter. We find that the imaginary part of the coupling plays a role analogous to a drift in diffusive systems. Furthermore, in the periodic chain case, the bias can be interpreted as a Peierls substitution, effectively incorporating the influence of a constant magnetic field into the hopping amplitudes. Traditionally, the study of quantum backflow is based on maximizing the time integral of the probability current over a finite time window, motivated by the fact that in one-dimensional systems this quantity is bounded. This approach produces a wave function characterized by a flux oscillating closely around a negative value throughout the time window of integration, yielding a large integrated flux, though its instantaneous intensity remains relatively low. For the tight-binding systems we study, a similar behavior occurs, however, we find that the maximum value of the integrated flux is larger for the tight-binding systems than for their continuous counterparts; for example, the integrated backflow for a particle on a continuous infinite line Bracken and Melloy (1994) is bounded by $c_{\text{BM}}=0.0384517$ compared to the value $0.0764734$ that can be reached in a discrete infinite chain. In the case of periodic chains, this bound depends on the number of sites, being larger for small lattice sizes and $\epsilon=0$. Going beyond this approach, in this paper, we investigate the maximum values that can be reached by the instantaneous backflow amplitude at a given time and position in discrete biased tight-binding chains. We find that, for both geometries considered, the maximum backflow appears as pronounced negative flux values that decay relatively rapidly in time. This behavior is particularly evident in the infinite-chain case. By contrast, for a periodic chain, large backflow values recur in time, although with less intensity than the maximum.
It is worth pointing out that the detection of backflow is still an experimental challenge, since it has not been observed in the laboratory, although some experimental proposals have been formulated Palmero et al. (2013); Yearsley et al. (2012); Barbier and Goussev (2021). While our approach is purely theoretical and the systems we study are extremely simple, the fact that they can sustain larger backflows than their continuous counterparts may broaden the possible experimental approaches in which this effect could be detected.
Finally, it would be interesting to extend our analysis to backflow and its effects in other systems, including, among others, one-dimensional SSH models Wei et al. (2025), quantum arrival-time distributions Dhar et al. (2015), Bloch wave packets in periodic potentials Goussev and Morozov (2025), disordered systems Cui et al. (2023), and quantum walks Zimborás et al. (2013).
However, as far as we know, in these models, even constructing the wave packets composed of non-negative momentum eigenstates necessary for studying quantum backflow, may pose a challenge.